The force on a charged particle due to electric and magnetic fields is given by F ⃗=qE ⃗+qv×B ⃗ . Suppose E ⃗ is along the X-axis and B ⃗ along the Y-axis. In what direction and with what minimum speed v should a positively charged particle be sent so that the net force on it is zero ?

Grade:11

1 Answers

Kevin Nash
9 years ago
18. Give F ⃗ qE ⃗ q(v ⃗ B ⃗) = 0 ⇒ E ⃗ (v ⃗ B ⃗) So, the direction of v ⃗ x B ⃗ should be opposite to the direction of E ⃗ . Hence, v ⃗ should be in the positive yz-plane. Again, E = vB sin θ ⇒ v = E/(B sin⁡θ ) For v to be minimum, θ = 90° and so Vmin = F/B So, the particle must be projected at a minimum speed of E/B along +ve z-axis (θ = 90°) as shown in the figure, so that the force is zero.